Suppose that the profit P obtained in selling x units of a certain item each week is

given by

p=50 √ 3 - 0.5-500(0<x<8000)

.


Find the rate of change of P with respect to x when x=1600.


1
Expert's answer
2021-10-05T13:49:42-0400
"p(x)=50\sqrt{3}x^2 - 0.5x-500"

"p'(x)=100\sqrt{3}x-0.5"

"p'(1600)=100\sqrt{3}(1600)-0.5"

"=160000\sqrt{3}-0.5"